# Negative Sampling in Recommendation: A Survey and Future Directions

### v1

@@ 3.5. Knowledge-enhanced Negative Sampling
As systematically demonstrated in Table 6, General KNS operates by exploiting the heterogeneous knowledge items. It outperforms the traditional negative sampling methods by utilizing the broader knowledge in multi-domain scenarios to sample the more informative negative samples with richer semantic meaning. It ensures that the selected negative samples are not only dissimilar to the positive samples but also incorporate meaningful semantic relationships with the positive samples, aligning more closely with the user’s preferences. However, it relies on high-quality knowledge data and additionally introduces data processing complexity. On the other hand, KG-based KNS leverages the structured information from the Knowledge Graph to enhance the quality of the sampled negative items. By incorporating the user-item interactions to the entities in KG, KG-based KNS is able to introduce the high-order structural coherence and the multi-hop node connectivity into negative samples, thereby precisely improving the semantic relevance rather than the feature relevance between positive and negative samples. However, it excessively depends on KG construction and faces challenges related to the computational complexity of modeling and maintaining the KG.
3.5.1. General KNS
The rich auxiliary relationships of users and the informative knowledge of items typically imply user preferences and item semantics, which could help overcome the sparsity issue in RS. General KNS selects negative samples based on the existing latent knowledge patterns to better understand the semantics and relations between users and items (Steck, 2018; Sun et al., 2015; Zhao et al., 2019; Krishnan et al., 2020). The learning process of the knowledge-based recommendation is grounded on contextual information. For instance, MMT-Net (Krishnan et al., 2020) designs the context-bias to reveal the novel tuples which display uncommon combinations of the interaction and its corresponding context, which can be defined as , where denotes the final context combinations. To enhance the training effectiveness of the proposed context-bias, MMT-Net develops a negative sampling approximation with two learning objectives to randomly identify the likely item given the user and interaction context and identify the likely context given the user and the item, respectively. Then the context-bias has been used to attenuate the loss for the tuple . Thus, the attenuated loss can be used to determine the embedding space and induce a novelty-weighted training curriculum focused on harder samples as training proceeds. SA_OCCF (Sun et al., 2015) leverages sentiment information extracted from user reviews in the dataset to assign the weight for each negative example rather than using a global weighting scheme in wAMAN (Pan et al., 2008). Instead of the review information, some works also leverage the informative genre and category knowledge for negative samplings. To list a few, LDA (Steck, 2018) devises a two-tiered sampling approach that first samples a genre from the learned genre-distribution for user , and then samples the movie as the negative sample from the learned distribution regarding the genre . IntentGC (Zhao et al., 2019) samples negative items in the same leaf category as the corresponding positive item to make the model capable of distinguishing hard items. Notably, if the sampled item exists in the interaction sequence, IntentGC discards it and re-samples a new negative item.
3.5.2. KG-based KNS
KG-based KNS leverages the explicit association and high-order correlation among users, items, and other entities in KG to sample negative instances (Liu et al., 2021b; Ali et al., 2020; Wang et al., 2020; Chen et al., 2022a). KG, which introduce extra relations among items and real-world entities (e.g., item attributes) with the comprehensive structural knowledge, are well known for its potential to infer informative negative samples and enhance recommenders’ accuracy and explainability. To precisely improve the semantic relevance between positive and negative samples and enhance the accessible information of the sampled negative instances, a series of works attempt to introduce the high-order structural coherence and the multi-hop node connectivity into negative samples in RS (Liu et al., 2021b; Ali et al., 2020; Wang et al., 2020). For example, AnchorKG (Liu et al., 2021b) design a simple but effective strategy to randomly select the in and conduct a random walk along a -hop path starting from the positive sample in KG. To minimize computational complexity in the model optimization, PR-HNE (Ali et al., 2020) employs unigram distribution to sample items (nodes) in candidates that appear in the KG independently of the presence of other items (nodes) over each interaction (edge). With the reinforcement learning (RL) paradigm, KGPolicy (Wang et al., 2020) samples knowledge-enhanced negative items from the relationships between items and entities in KG, emphasizing KG entities and items linked within two-hop paths and facilitating the effective learning procedure of the recommender. KGAttack (Chen et al., 2022a) first proposes knowledge-enhanced candidate selection to employ KG for localizing some relevant item candidates with similar attributes. Subsequently, it leverages RL paradigm to calculate the sampling probabilities for each item in the candidates by calculating the similarity state between the current state representations and graph embeddings.
@@ 3.5.2. KG-based KNS
| Cross-domain Recommendation | MMT-Net (Krishnan et al., 2020) | Knowledge-based IRW | SIGIR | 2020 |
@@ 4. Negative Sampling in Multiple Practical Recommendation Scenarios
The field of recommendation systems is replete with a vast spectrum of diverse recommendation scenarios. This diversity stems from user demands, historical behaviors, and heterogeneous applications, thereby incorporating the field with a blend of challenges and prospects (Gao et al., 2022b; Chen et al., 2023a). Benefits from its superior performance and easy-to-employ character, negative sampling methods have garnered significant attention from researchers and practitioners in recommendation (Zhang et al., 2013; Shi et al., 2023; Ding et al., 2019; Lian et al., 2020a). Some researchers have attempted to design a universal approach to constructing a negative sampling method that is suitable for all recommendation scenarios (Ma et al., 2023). Nonetheless, the employment, objective, and data availability significantly vary across diverse recommendation scenarios, leading to the absence of a negative sampling algorithm that can comprehensively encompass all recommendation scenarios. In this section, we undertake an in-depth exploration of the distinctiveness of various recommendation scenarios, delineating subtle disparities in the negative sampling methodologies under the same recommendation scenarios. Ultimately, we conclude with a comprehensive discussion and outlook concerning the paradigm of tailored negative sampling techniques for each specific scenario. Specifically, the subsequent discussions span five pivotal recommendation scenarios in the following subsections: Collaborative-guided Recommendation, Sequential Recommendation, Multi-modal Recommendation, Multi-behavior Recommendation, Cross-domain Recommendation and CL-enhanced Recommendation. These discrete discussions and analyses illustrate the adaptability and customization inherent to negative sampling strategies, a fundamental prerequisite for optimizing recommendation algorithms in diverse recommendation contexts.
4.1. Introduction of Negative Feedback in Practical Recommendation
Both positive and negative feedback can reflect the user’s preferences on diverse items in the corpus, which are equally essential for modeling user preferences in RS (Zhao et al., 2018; Wu et al., 2022b). Apart from explicit feedback and implicit feedback, the presence of real negative feedback is a notable factor in practical RS (Xie et al., 2021). Leveraging such data to its maximum capacity can provide the recommender with a significant information gain. Constrained by the exposure bias and popularity bias inherent in real-world RS, the interaction matrix reveals a notable sparsity, which constitutes a computationally insurmountable problem in RS (Huang et al., 2023). As mentioned in Sec. 2.1, negative sampling emerge as a critical and irreplaceable element in recommendation. How to effectively discover users’ real negative interests in negative sampling stands as a central research objective (Chen et al., 2023c), the detailed expressions of the associated concepts and methodologies are described in Sec. 3. Within this section, our principal endeavor is to introduce the two types of feedback mentioned above 777Due to the scarcity of real user behaviors, we discuss their definition and application in Sec .4.5..
Explicit Feedback: Explicit Feedback materializes in the form of precise ratings to provides unequivocal insights into user preferences (e.g., from 1.0 to 5.0). Actually, this may not necessarily reflect the users’ real preferences. Even for the item explicitly rated as 1.0, they have been chosen by this user from an extensive candidate pool. Therefore, such negative feedback reflects user preferences to a certain extent, that is, they may not be as too ”negative”.
Implicit Feedback: Implicit Feedback encompasses data derived from user behaviors without any ratings and reviews (e.g., residence time and whether to choose dislike). It lacks explicitness and necessitates the model-based inferences to elucidate users’ preferences, which may introduce long-tailed information and exposure bias in RS.
@@ 4.1. Introduction of Negative Feedback in Practical Recommendation
denotes “unobserved samples”.From the above two types of feedback, implicit feedback in the observations is fundamentally different from explicit feedback, which is captured in a structured manner (Saito et al., 2020). Precisely, explicit feedback delivers direct insights into user preferences, while implicit feedback unravels concealed user interests. Apart from the implicit and explicit feedback, real-world RS typically involves the concept of real negative feedback, which hides in real feedback and contributes to a more holistic understanding of user preference (Xie et al., 2021; Huang et al., 2023). It refers to the authentic user negative behaviors, with its definition intimately linked to various recommendation scenarios. For instance, non-purchases are considered negative samples in the context of purchases, while in the context of clicks, exposures without clicks constitute negative samples. Similarly, non-likes, in comparison to likes, are treated as negative samples. This encapsulates explicit information of users’ disapproval and unclicks, contributing to a more comprehensive understanding of user interest (Ding et al., 2023). These authentic feedback wields significant influence in recommendation. With the specific recommendation scenario, negative sampling algorithms can be designed and optimized according to the distinct types of feedback data (See Sec.4.5).
4.2. Collaborative-guided Recommendation
Collaborative-guided Recommendation (CR) aims to gather and analyze all the collaborative behaviors for predicting the potential preference of a target user (Rendle et al., 2012). It typically focuses on projecting users and items into the latent embedding space to reflect users’ preferences (He et al., 2017). In contrast to multi-modal recommendation and cross-domain recommendation, the available data of CR is exclusively restricted in the user-item interactions, which constrains its practical accuracy. Moreover, it is widely recognized that merely considering positive instances is insufficient to capture the intricacies of user preferences (Zhang et al., 2013; Shi et al., 2023). In light of this, a series of CF methods have investigated how to maximize the utilization of interaction data for the purpose of sampling more informative negative instances in CF tasks (Chen et al., 2020c; Shi et al., 2023; Ding et al., 2020; Wu et al., 2022a).
Constrained by the data availability inherent to CF tasks, prevailing negative sampling methods typically employ SNS approaches to facilitate the process of negative sample selection. Precisely, randomly sample items from the unobserved corpus with the same probability (He et al., 2023a, 2017) are extensively employed in CF. A substantial number of CF approaches also focus on the dynamic selection of negative samples, which are information-rich and contextually suitable in a current context during the training of CF models. DNS (Zhang et al., 2013) and DNS (Shi et al., 2023) regard the items with the highest score as the most informative samples and attempt to select them as the negative samples. SRNS (Ding et al., 2020), GDNS (Zhu et al., 2022), and DENS (Lai et al., 2023) leverage the inherent distribution (e.g., statistical information and disentangled representation) between positive and negative samples to dynamically select the suitable negative instances under the current context in training. Furthermore, a CL-based CF algorithm EGLN (Yang et al., 2021d) attempts to conduct the structural augmentation on the interaction graph to maximize the local-global consistency in collaborative representation learning.
4.3. Sequential Recommendation
Sequential Recommendation (SR) is a straightforward and effective approach for capturing users’ dynamic preference patterns from their historical behaviors (Dallmann et al., 2021; Wang et al., 2023a). Its goal is to predict the next items the user is likely to interested with by modeling the temporal correlations in user behavioral sequence. Some studies have proven the effectiveness of hard negative sampling strategies in SR (Wu et al., 2019b; Chen et al., 2022b; Xie et al., 2022a). Through the selection of negative examples that exhibit a certain similarity to positive samples, these strategies bolster the recommender’s capability to finely distinguish users’ dynamic preferences in recommendation. Most of the SR methods randomly select unobserved items from the corpus as the negative samples (Yang et al., 2022b; Dallmann et al., 2021; Petrov and Macdonald, 2022). With the development of contrastive learning, some studies rely on diverse sequence augmentation methods (e.g., crop, mask, reorder, substitute, and insert) to generate negative samples in CL tasks, assisting in modeling users’ temporal preferences. Despite the utilization of the hard negative sampling strategy, DEEMS (Wu et al., 2019b) and GeoSAN (Lian et al., 2020b) have not considered the temporal relationships among behaviors, opting to treat each interaction with equitable importance. Similarly, ELECRec (Chen et al., 2022b) attempts to leverage the discriminator in GAN to distinguish the target sample, overlooking the sequential knowledge within the behavioral sequence. Evidently, negative sampling holds particular significance for SR, which is a concern that existing research has momentarily disregarded and may serve as a promising avenue of exploration for future scholars.
4.4. Multi-modal Recommendation
Multi-modal Recommendation (MMR) is crucial in real-world recommendation applications with massive modality-specific information, that is, micro-videos, images, audio, and text (Liu et al., 2019; Tao et al., 2022). Since the multi-modal information characteristics may reflect users’ actual preferences within the fine-grained modality level (Chen et al., 2023b), the idealized multi-modal recommender possesses the capacity to uncover the hidden relationships between different modalities and recover the complementary information that can not be captured by the traditional CF methods (Dong et al., 2024). To achieve satisfactory performance, MMR methods typically require a substantial amount of both positive (present in the observed interactions) and negative items (obscured within the unobserved interactions) to train the recommenders in a supervised manner. This motivates researchers to explore how to sample more informative negative instances that better align with users’ actual preferences with the multi-modal knowledge. Similar to other recommendation scenarios, most of related studies commonly apply SNS (uniform SNS (Liu et al., 2019; Wu et al., 2019a) and predefined SNS (Togashi et al., 2021; Wang et al., 2019b)) to select negative samples in MMR. Some works also investigate the hidden correlations within the items’ multiple modalities to conduct the modalities-aware CL pairs. For the traditional negative sampling strategies, FairNeg (Chen et al., 2023b) achieves debiased recommendation via capturing the group-level unfairness and balancing the sampling probabilities of each semantic group. RPRM (Wang et al., 2021d) attempts to model the review properties’ consistency among different items and consequently sample the negatives with similar review properties of positive samples. POWERec (Dong et al., 2024) leverages the strength of the multi-modal representations learned from the identical item to enhance user interest understanding in the unreliable modality, making all the modality-specific user interests to be fully learned. These negative sampling strategies can incorporate the multi-modal information to intelligently refine the selection of the relevant negative instances, which exhibits meaningful semantic relationships with the positive samples and closely aligns with the user’s intricate interests.
@@ 4.5. Multi-behavior Recommendation
Benefits from the multiple types of interaction between users and items in real-world scenarios, multi-behavior recommendation (MBR) refers to taking full advantage of these heterogeneous user behaviors (e.g., skipped, clicked, and ordered) to alleviate the data sparsity or cold-start issues (Xie et al., 2021; Xia et al., 2022; Gu et al., 2021). Its core idea is to model the cross-type behavior dependency in real-world RS, which can provide complementary information for user’s interests encoding (Gu et al., 2021; Zhou et al., 2018; Yang et al., 2021a). In such an application, effectively modeling the multi-type user behaviors can provide auxiliary knowledge and precise user preference to characterize the underlying semantics of the user’s actual interest representation (Zheng et al., 2018; Gu et al., 2022). Existing MBR works primarily consider the multi-behavior dependencies with the predefined correlations (Yang et al., 2022b; Xia et al., 2022), while seldom embarking on the selection of diverse negative samples for individual behaviors through negative sampling. To some extent, this limits its capacity to capture the intricate cross-type behavior dependencies within complex real-world recommendation scenarios. For instance, ZEUS (Gu et al., 2021) and HMG-CR (Yang et al., 2021a) independently construct negative CL pairs by forming inter-connections between multiple behaviors and generating behavior-aware hyper meta-graphs. It enables the simultaneous modeling of the diverse interaction patterns and underlying cross-type behavior inter-dependencies, verifying the positive effects of incorporating the cross-type of multi-behavioral context in MBR.

### v2

@@ 3.5. Knowledge-enhanced Negative Sampling
3.5.1. General KNS
The rich auxiliary relationships of users and the informative knowledge of items typically imply user preferences and item semantics, which could help overcome the sparsity issue in RS. General KNS selects negative samples based on the existing latent knowledge patterns to better understand the semantics and relations between users and items (Steck, 2018; Sun et al., 2015; Zhao et al., 2019; Krishnan et al., 2020). The learning process of the knowledge-based recommendation is grounded on contextual information. For instance, MMT-Net (Krishnan et al., 2020) designs the context-bias to reveal the novel tuples which display uncommon combinations of the interaction and its corresponding context, which can be defined as , where denotes the final context combinations. To enhance the training effectiveness of the proposed context-bias, MMT-Net develops a negative sampling approximation with two learning objectives to randomly identify the likely item given the user and interaction context and identify the likely context given the user and the item, respectively. Then the context-bias has been used to attenuate the loss as , thereby inducing a novelty-weighted training curriculum focused on harder samples as training proceeds. SA_OCCF (Sun et al., 2015) leverages sentiment information extracted from user reviews to assign the weight for each negative example rather than the global weighting scheme. Instead of the review information, some works also leverage the informative genre and category knowledge for negative samplings. To list a few, LDA (Steck, 2018) devises a two-tiered sampling approach that first samples a genre from the learned genre-distribution for user , and then samples the movie as the negative sample from the learned distribution regarding the genre . IntentGC (Zhao et al., 2019) samples negative items in the same leaf category as the corresponding positive item to make the model capable of distinguishing hard items.
3.5.2. KG-based KNS
KG-based KNS leverages the explicit association and high-order correlation among users, items, and other entities in KG to sample negative instances (Liu et al., 2021b; Ali et al., 2020; Wang et al., 2020; Chen et al., 2022a). KG, which introduces extra relations among items and real-world entities (e.g., item attributes) with the comprehensive structural knowledge, are well known for its potential to enhance recommenders’ accuracy and explainability. To precisely improve the semantic relevance between positive and negative samples and enhance the accessible information of the sampled negative instances, a series of works attempt to introduce the high-order structural coherence and the multi-hop node connectivity into negative sampling in RS (Liu et al., 2021b; Ali et al., 2020; Wang et al., 2020). For example, AnchorKG (Liu et al., 2021b) design a simple but effective strategy to randomly select the in and conduct a random walk along a -hop path starting from the positive sample in KG. To minimize computational complexity, PR-HNE (Ali et al., 2020) employs unigram distribution to sample items (nodes) in candidates that appear in the KG independently of the presence of other items over each interaction (edge). With the reinforcement learning (RL) paradigm, KGPolicy (Wang et al., 2020) samples knowledge-enhanced negative items from the relationships between items and entities in KG, emphasizing KG entities and items linked within two-hop paths and facilitating recommenders’ learning process. KGAttack (Chen et al., 2022a) proposes knowledge-enhanced candidate selection to employ KG for localizing some relevant item candidates with similar attributes. Subsequently, it leverages RL paradigm to obtain the sampling probabilities of each candidate item by calculating the similarity between the current state representations and graph embeddings.
@@ 3.5.2. KG-based KNS
Summary: As demonstrated in Table 6, General KNS exploits all available heterogeneous knowledge to sample the more informative negative samples with richer semantic meaning, thereby outperforming the traditional negative sampling methods. It ensures that the selected negative samples are not only dissimilar to the positive samples but also incorporate meaningful semantic relationships with the positive samples, aligning more closely with users’ preferences. However, it relies on high-quality knowledge and additionally introduces data processing complexity. In contrast, KG-based KNS leverages the structured information from KG to enhance the quality of the sampled negatives. By incorporating the user-item interactions to the entities in KG, KG-based KNS can introduce the high-order structural coherence and the multi-hop node connectivity into negative samples, thereby precisely improving the semantic relevance rather than the feature relevance between positive and negative samples. However, it excessively depends on KG construction and faces challenges related to the computational complexity of modeling and maintaining the KG.
The field of recommender systems is replete with a vast spectrum of diverse scenarios, which stems from user demands, historical behaviors, and heterogeneous applications, thereby incorporating the field with a blend of challenges and prospects (Gao et al., 2024; Chen et al., 2023a). Benefits from its superior performance and easy-to-employ character, negative sampling has garnered significant attention from researchers and practitioners in recommendation (Zhang et al., 2013; Shi et al., 2023; Ding et al., 2019; Lian et al., 2020a). Some researchers have attempted to design a universal negative sampling method that is suitable for all recommendation scenarios (Ma et al., 2023). Nonetheless, the employment, objective, and data availability significantly vary across diverse scenarios, leading to the absence of a negative sampling algorithm that can comprehensively encompass all recommendation scenarios. In this section, we undertake an in-depth exploration of the distinctiveness of various recommendation scenarios, delineating subtle disparities in the negative sampling methodologies under the same scenarios. Ultimately, we conclude with a comprehensive discussion and outlook concerning the paradigm of tailored negative sampling techniques for each specific scenario. Specifically, the subsequent discussions span six pivotal recommendation scenarios in the following subsections: Collaborative-guided Recommendation, Sequential Recommendation, Multi-modal Recommendation, Multi-behavior Recommendation, Cross-domain Recommendation and CL-enhanced Recommendation. These discrete discussions and analyses illustrate the adaptability and customization inherent to negative sampling strategies, a fundamental prerequisite for optimizing recommendation algorithms in diverse recommendation contexts.
@@ 4. Negative Sampling in Multiple Practical Recommendation Scenarios
| Cross-domain Recommendation | MMT-Net (Krishnan et al., 2020) | Knowledge-aware IRW | SIGIR | 2020 |
@@
4.1. Collaborative-guided Recommendation
Collaborative-guided Recommendation (CR) aims to gather and analyze all the collaborative behaviors for predicting users’ potential preferences (Rendle et al., 2009). It typically focuses on projecting users and items into the latent embedding space to reflect their correlation (He et al., 2017). In contrast to other recommendation scenarios, the available data of CR is exclusively restricted in the user-item interactions, which constrains its practical accuracy. Moreover, it is widely recognized that merely considering positive instances is insufficient to capture the intricacies of user preferences (Zhang et al., 2013; Shi et al., 2023). In light of this, a series of CF methods have investigated how to maximize the utilization of interaction data to sample more informative negative instances (Chen et al., 2020c; Shi et al., 2023; Ding et al., 2020; Wu et al., 2022a).
Constrained by the data availability inherent to CR tasks, prevailing negative sampling methods typically employ SNS approaches to facilitate the process of negative sample selection. Precisely, randomly sampling items from the unobserved corpus with the same probability (He et al., 2023a, 2017) are extensively employed in CR. A substantial number of CR approaches also focus on the dynamic selection of negative samples, which are information-rich during the training of CR models. DNS (Zhang et al., 2013) and DNS (Shi et al., 2023) regard the items with the highest score as the most informative samples and attempt to select them as the negative samples. SRNS (Ding et al., 2020), GDNS (Zhu et al., 2022), and DENS (Lai et al., 2023) leverage the inherent distribution (e.g., statistical information and disentangled representation) between positive and negative samples to dynamically select the suitable negative instances under the current context in training. Furthermore, a CL-based CR algorithm, EGLN (Yang et al., 2021d), attempts to conduct the structural augmentation on the interaction graph to maximize the local-global consistency in collaborative representation learning.
4.2. Sequential Recommendation
Sequential Recommendation (SR) is a straightforward but effective approach for capturing users’ dynamic preference patterns (Dallmann et al., 2021; Wang et al., 2023a; Ma et al., 2024c). Its goal is to predict the next items the user is likely to interacted with by modeling the temporal correlations in his/her behavioral sequence. Some studies have proven the effectiveness of hard negative sampling strategies in SR (Wu et al., 2019b; Chen et al., 2022b; Xie et al., 2022a). Through the selection of negative examples that exhibit a certain similarity to positive samples, these strategies bolster the recommender’s capability to finely distinguish users’ dynamic preferences. Most existing SR methods randomly select unobserved items from the corpus as the negative samples (Yang et al., 2022b; Dallmann et al., 2021; Petrov and Macdonald, 2022). With the development of contrastive learning (CL), some studies rely on diverse sequence augmentation methods (e.g., crop, mask, reorder, substitute, and insert) to generate negative samples, assisting in modeling users’ temporal preferences. Despite the utilization of hard negative sampling strategies, DEEMS (Wu et al., 2019b) and GeoSAN (Lian et al., 2020b) opt to treat each interaction with equitable importance rather than explicitly considering the temporal relationships among behaviors Similarly, ELECRec (Chen et al., 2022b) leverages discriminator in GAN to distinguish the target sample, overlooking the sequential knowledge within behavioral sequences. Evidently, negative sampling holds particular significance for SR, which has been momentarily disregarded by existing researchers and may serve as a promising avenue of exploration for future scholars.
4.3. Multi-modal Recommendation
Multi-modal Recommendation (MMR) is crucial in real-world recommendation applications with massive modality-specific information, that is, micro-videos, images, audio, and text (Wang et al., 2023a; Tao et al., 2022; Ma et al., 2024d). Since the multi-modal information may reflect users’ actual preferences within the fine-grained modality level (Chen et al., 2023b), the idealized multi-modal recommender possesses the capacity to uncover the hidden relationships between different modalities and recover the complementary information that can not be captured by CR methods (Dong et al., 2024). To achieve satisfactory performance, MMR methods typically require a substantial amount of both positive (present in the observed interactions) and negative items (obscured within the unobserved interactions) to train recommenders in a supervised manner. This motivates researchers to explore how to sample more informative negative instances that better align with users’ actual preferences with the multi-modal knowledge. Similar to other recommendation scenarios, most related studies commonly apply SNS (Wang et al., 2023a; Wu et al., 2019a; Wang et al., 2019) to select negative samples in MMR. Some works also investigate the hidden correlations within the items’ multiple modalities to conduct the modality-aware CL pairs. For the traditional negative sampling strategies, FairNeg (Chen et al., 2023b) achieves debiased recommendation via capturing the group-level unfairness and balancing the sampling probabilities of each semantic group. RPRM (Wang et al., 2021d) attempts to model the review properties’ consistency among different items and consequently sample the negatives with similar review properties of positive samples. POWERec (Dong et al., 2024) leverages the strength of multi-modal representations learned from the identical item to enhance user interest understanding in the unreliable modality, making all the modality-specific user interests to be fully learned. These algorithms can incorporate the multi-modal information to intelligently refine the selection of the relevant negative instances, which exhibits meaningful semantic relationships with the positive samples and closely aligns with the user’s intricate interests.
@@ 4.3. Multi-modal Recommendation
denotes “unobserved samples”.
@@ 4.4. Multi-behavior Recommendation
As illustrated in Fig. 8, real-world RS typically involves the concept of real negative feedback in addition to implicit and explicit feedback, which contributes to holistic understanding of user preferences (Xie et al., 2021; Huang et al., 2023). It refers to the authentic user negative behaviors, with its definition intimately linked to various recommendation scenarios. For instance, non-purchases are considered negative samples in the context of purchases, while in the context of clicks, exposures without clicks constitute negative samples. Similarly, non-likes, in comparison to likes, are treated as negative samples. This encapsulates explicit information of users’ disapproval, wielding significant influence in recommendation (Ding et al., 2023).
Benefits from the multiple types of interaction between users and items in real-world scenarios, multi-behavior recommendation (MBR) refers to taking full advantage of these heterogeneous user behaviors (e.g., skipped, clicked, and ordered) to alleviate the data sparsity or cold-start issues (Xie et al., 2021; Xia et al., 2022; Gu et al., 2021; Zhang et al., 2024a). As illustrated in Fig. 9, its core idea is to model the cross-type behavior dependency in RS, which can provide complementary information for user’s interest encoding (Gu et al., 2021; Zhou et al., 2018; Yang et al., 2021a). In such an application, effectively modeling the multi-type user behaviors can provide auxiliary knowledge and precise user preference to characterize the underlying semantics of user’s actual interest representation (Zheng et al., 2018; Gu et al., 2022). Existing MBR works primarily consider the multi-behavior dependencies with the predefined correlations (Yang et al., 2022b; Xia et al., 2022), while seldom embarking on the selection of diverse negative samples for individual behaviors through negative sampling. To some extent, this limits its ability to capture the intricate cross-type behavior dependencies within complex real-world RS scenarios. For instance, ZEUS (Gu et al., 2021) and HMG-CR (Yang et al., 2021a) independently construct negative CL pairs by forming inter-connections between multiple behaviors and generating behavior-aware hyper meta-graphs. It enables the simultaneous modeling of the diverse interaction patterns and underlying cross-type behavior inter-dependencies, verifying the positive effects of incorporating the cross-type of multi-behavioral context in MBR.